Total mass 57.3 kg. Rides .64km in 14.6 min in constant speed. Aceleración is 9.8 m/s2 what is the kinetic energy?


1
Expert's answer
2022-02-09T13:39:37-0500

Explanation & calculations


  • Since the whole motion happens at a constant speed, there is no room for an acceleration parallel to the body's direction of motion.
  • Therefore, the given acceleration should be perpendicular to the body's direction of motion.


  • Then the constant velocity is

"\qquad\qquad
\begin{aligned}
\small v&\small \frac{0.64\times1000m}{14.6\times60\,s}\\
&=\small 0.73\,ms^{-1}
\end{aligned}"

  • since the speed is constant, the kinetic energy is constant and that can be calculated by the formula "\small E= \frac{1}{2}mv^2"